# Fanxiao Li 1, Jiaying Wu2, Tingchao Fu1, Yunyun Dong1, Bingbing Song1, Wei Zhou 1

### v1

@@ Abstract
Content Warning: this paper contains potentially misleading text and images.
1 Introduction
Multimodal misinformation, often in the form of image-text combinations, poses escalating threats to public discourse, societal trust, and civic stability (Bovet and Makse 2019; Murayama 2021; Wang et al. 2024a; Wu et al. 2025b). To counter this threat, Large Vision-Language Models (LVLMs) (Hurst et al. 2024; Anil et al. 2023; Liu et al. 2024a; Meta 2024) has emerged as a dominant approach for multimodal misinformation detection (MMD), offering strong multimodal reasoning and retrieval-augmented verification capabilities (Kangur et al. 2025; Qi et al. 2024; Braun et al. 2024; Li et al. 2025a; Geng et al. 2024; Wan et al. 2025).
@@ 1 Introduction
However, the rise of generative AI (GenAI) systems such as GPT-4o (Hurst et al. 2024) is reshaping the information landscape. GenAI enables the creation of highly diverse and stylistically varied content at scale (Matich, Thomson, and Thomas 2025; Kiskola et al. 2025; Davenport and Mittal 2023; Nishal and Diakopoulos 2024; Kieslich, Diakopoulos, and Helberger 2024), fundamentally transforming how news is composed, visualized, and disseminated. We refer to this phenomenon as GenAI-driven News Diversity, which manifests in two key forms: (1) Controlled News Diversity, where users produce semantically consistent variants of the same news content through rewording, paraphrasing, or shifts in visual framing; and (2) Open-ended News Diversity, where GenAI generates entirely novel content via text-to-image synthesis or narrative reconstruction.
While this diversity enhances expressiveness and audience reach, it also introduces substantial challenges for LVLM-based MMD systems. Most existing models either assess image-text coherence using internal knowledge (Wang et al. 2024b) or retrieve external evidence for fact verification (Li et al. 2025b; Qi et al. 2024). Both approaches become fragile under content variation, which gives rise to what we term multi-level drift. As illustrated in Figure 1, we identify two key drift phenomena: (1) Model-Level Misperception Drift: Variations in surface form, including lexical, syntactic, or visual aspects, may distort the model’s perception, even when semantics remain unchanged. This issue is exacerbated by LVLMs’ known hallucination tendencies (Bai et al. 2024; Liu et al. 2024b). (2) Evidence-Level Drift: Content diversity compromises evidence retrieval by reducing the likelihood of finding exact matches, leading to loosely relevant or misleading external evidence (Chung et al. 2023; Xiao and Wang 2021; Feng et al. 2020). Beyond natural drift, GenAI also enables adversarial manipulation through guided retrieval attacks, wherein malicious actors flood the web with strategically crafted synthetic content. Such contamination can bias retrieval pipelines, poisoning the fact-checking process and degrading LVLM performance (Xu et al. 2024; Dai et al. 2023a). Despite their growing risk, these vulnerabilities remain insufficiently studied.
@@
To this end, we introduce DriftBench, a large-scale benchmark consisting of 16,000 news instances spanning six diversified categories across both image and text modalities, with human validation to ensure semantic fidelity. Specifically, we source high-quality real and fake samples from NewsCLIPpings (Luo and Trevor Darrell 2021) and systematically diversify them using an automated GenAI-based pipeline grounded in our news diversity taxonomy. For real news, we apply controlled diversification to the image, text, or both while preserving factual consistency. For fake news, we incorporate both controlled and open-ended transformations, rewriting text to fabricate false narratives and generating misleading images aligned with the fabricated content. The resulting dataset spans eight structured content categories, enabling rigorous evaluation of LVLM robustness under GenAI-driven variation.
With DriftBench, we design three corresponding evaluation tasks: (1) Performance under Diversity: Evaluate how SOTA LVLM-based detectors handle controlled and open-ended content variation, focusing on degradation due to model- and evidence-level drift. (2) Robustness to Malicious Evidence Contamination: Simulate guided retrieval attacks by injecting GenAI-generated misleading evidence into the verification pipeline. (3) Reasoning Behavior Analysis: Employ an LVLM-as-a-Judge framework (Li et al. 2024; Chen et al. 2024; Pu et al. 2025) to assess models’ explanation quality and attribution patterns under diverse input conditions.
@@ 2 Related Work
NewsCLIPpings (Luo and Trevor Darrell 2021) OOC misinformation - - -
VERITE (Papadopoulos et al. 2024) Real-world misinformation - - -
MMFakeBench (Liu et al. 2024c) Multi-source misinformation - - -
MFC-Bench (Wang et al. 2024b) Multimodal fact-checking - - - -
MiRAGenNews (Huang et al. 2024) AI-generated misinformation - - -
@@
2.1 LVLM-Based MMD Approaches
@@
LVLM-based multimodal fact-checking is the current mainstream method for multimodal misinformation detection. SNIFFER (Qi et al. 2024) performs two-stage fine-tuning based on InstructBLIP (Dai et al. 2023b), aiming to detect misinformation from the perspective of entity matching and evidence verification. CMIE (Li et al. 2025b) and E2LVLM (Wu et al. 2025a) further explore how to discover deeper relationships and optimize the selection and usage of external evidence. LRQ-FACT (Beigi et al. 2024) and LEMMA (Xuan et al. 2024) optimize retrieval strategies to acquire more targeted external evidence. These methods fully investigate the capabilities of LVLMs in retrieval, comprehension, and reasoning, and are optimized around the acquisition and utilization of high-quality evidence. However, most of them assume that the input claim is static and expressed in a single form, lacking systematic investigation and evaluation of such methods’ performance under news diversity and the resulting multi-level drift.
2.2 Multimodal Misinformation Benchmarks
@@
To evaluate the performance of MMD detectors, several benchmarks have been proposed. DGM4 (Shao, Wu, and Liu 2023) targets fine-grained manipulation of visual and textual modalities, suitable for manipulated misinformation detection. NewsCLIPpings (Luo and Trevor Darrell 2021) presents OOC misinformation by constructing mismatched image-text pairs with external evidence. VERITE (Papadopoulos et al. 2024) collects real-world OOC misinformation cases from fact-checking platforms, exploring modality bias and its impact. MiRAGenNews (Huang et al. 2024) uses GenAI techniques to generate image-text pairs of misinformation, evaluating AI-generated misinformation detection. MMFakeBench (Liu et al. 2024c) aggregates misinformation from both real-world and GenAI sources, assessing model robustness under mixed-source conditions. MFC-Bench (Wang et al. 2024b) probes the internal fact-checking abilities of LVLMs.
@@
In contrast, our proposed DriftBench is the first benchmark to systematically investigate the challenges introduced by GenAI-driven news diversity, with a particular focus on two underexplored phenomena: model misperception drift and evidence drift. By introducing semantically diverse real and fabricated variants through controlled and open-ended transformations, DriftBench fills a critical gap in existing benchmarks. A comparative overview is provided in Table 1.
3 DriftBench
@@
To systematically assess the impact of GenAI-driven news diversity and the resulting multi-level drift on LVLM-based multimodal misinformation detection, we introduce DriftBench, a large-scale benchmark comprising 16,000 news instances across eight categories. These categories are defined by two base types (real and fake news, each consisting of human-written text and authentic images) and six diversification strategies. DriftBench is constructed using an automated GenAI-based synthesis pipeline applied to real-world news content, enabling scalable, controlled, and semantically faithful diversification. Each instance consists of an image-text news pair, either original or diversified, accompanied by external evidence retrieved from the web. An overview of the dataset construction and evaluation framework is illustrated in Figure 2.
@@
3.1 Taxonomy of GenAI-Driven News Diversity
@@
  • F_OI_OT: An out-of-context (OOC) misinformation instance (Luo and Trevor Darrell 2021; Qi et al. 2024), created by mismatching the image and text from two different but individually trustworthy news reports—misleading when paired together.
@@
3.2 Dataset Creation
Based on the proposed taxonomy, we construct DriftBench, a large-scale benchmark comprising real and fake news instances systematically diversified through GenAI-driven content variation. We begin by sourcing 4,000 high-quality samples (2,000 real and 2,000 fake) from the NewsCLIPpings dataset (Luo and Trevor Darrell 2021). Real instances comprise semantically aligned image-caption pairs published by reputable news outlets. Fake instances are constructed via OOC manipulation, where the image and text are drawn from two different trustworthy reports. While individually accurate, their combination results in a misleading narrative with high surface plausibility.
@@ 3.2 Dataset Creation
Text diversifications and fabrications are generated using GPT-4o (Hurst et al. 2024). For text diversification (DT), we prompt the model to rephrase the original text while preserving its semantics. For text fabrication (FT), we prompt the model to generate a false narrative by altering the described event. Image diversifications (DI) are generated using FLUX.1 Redux [dev] (an open-source image variant generation model) (Labs et al. 2025), the model generates visually varied but semantically consistent depictions of the original content. Image fabrications (FI) are generated using FLUX.1 [dev] (an open-source text-to-image model) (Labs 2024), we condition the image generation on the fabricated narrative FT, yielding fully synthetic misinformation instances.
@@
For each image-text pair, we retrieve external evidence to support LVLM-based multimodal fact-checking. For image-based retrieval, we use a web crawler to identify webpages containing the image and extract their page titles as evidence. For text-based retrieval, we adopt the query generation strategy from LEMMA (Xuan et al. 2024), using the Serper API [^1] https://serper.dev/ to retrieve relevant web documents.
@@
3.3 Data Quality Assessment
@@
We report two metrics to assess data quality: (1) semantic preservation accuracy, computed from aggregated human annotations, and (2) inter-annotator agreement using Fleiss’ $\kappa$ (Fleiss 1971). Results show that DriftBench exhibits high semantic consistency and labeling reliability, with 98.6% accuracy and a Fleiss’ $\kappa$ of 0.872 for real-instance variants, and 96.6% accuracy with a Fleiss’ $\kappa$ of 0.741 for controlled fake variants.
3.4 Evaluation Tasks and Metrics
@@
Following established practices in LVLM-based MMD evaluation (Li et al. 2025b; Xuan et al. 2024; Qi et al. 2024), we report Accuracy and F1 as the primary metrics for detection performance.
4 Experiments
4.1 Experimental Setup
@@
Vanilla LVLMs. We assess three widely used models: (1) GPT-4o-mini (Hurst et al. 2024): a proprietary large vision-language model. (2) Claude-3.7-Sonnet (Anthropic 2025): a large reasoning model with multimodal capabilities. (3) Qwen-VL-Plus (Bai et al. 2023): a strong open-source vision-language model. Prompting strategies and instruction templates used for Vanilla LVLMs are detailed in Appendix (Figure 11).
LVLM-Based MMD Methods. We also evaluate three state-of-the-art MMD approaches. (1) SNIFFER (Qi et al. 2024), which fine-tunes LVLMs and retrieves external evidence directly through image-based querying. (2) CMIE (Li et al. 2025b), which models deep semantic alignment between image and text while mitigating noise in retrieved evidence. and (3) LEMMA (Xuan et al. 2024), which employs a modular pipeline that first generates optimized queries and then retrieves high-quality evidence.
4.2 Performance under GenAI-Driven Diversity
@@
4.3 Robustness to Evidence Contamination
Contaminated evidence further degrades model performance and can steer predictions in misleading directions. As discussed in Section 3.4, malicious actors can manipulate retrieval pipelines by injecting fabricated or misleading content into the evidence set. Table 3 shows that introducing such adversarial evidence causes a clear performance drop across all evaluated methods. Detailed results in Appendix A.6 indicate that real instances are increasingly misclassified as fake when contaminated evidence is present. For fake instances, the misleading evidence often supports the false narrative, leading models to mistakenly classify them as real and thus producing the outcome intended by the adversary.
@@ 4.3 Robustness to Evidence Contamination
Table 3: Comparison of LVLM-based multimodal misinformation detection performance under malicious evidence contamination, evaluated on the DI_OT setting within controlled news diversity.
@@
4.4 Model Reasoning Behavior Analysis
Figure 5: Reasoning explanations generated by LVLMs under controlled settings (R_OI_OT, R_DI_OT), open ended scenarios (F_FI_FT), and cases involving malicious evidence contamination (F_DI_OT). Definitions of each instance type are provided in Section 3.1.
@@ 4.4 Model Reasoning Behavior Analysis
Figure 6: Error attribution analysis showing the distribution of error types across different methods for real and fake instances.
@@ 4.5 Case Study
5 Conclusion
We present DriftBench, the first benchmark explicitly designed to evaluate the vulnerabilities of LVLM-based misinformation detection in the face of GenAI-driven news diversity. Our large-scale, systematically diversified dataset demonstrates that both controlled and open-ended content variation, along with adversarial evidence contamination, can substantially weaken the performance of current LVLM-based MMD systems. Beyond accuracy degradation, our analysis reveals distinct reasoning behaviors and error patterns among different model classes, reflecting varied reliance on visual-textual coherence and external evidence. These findings emphasize the need for robust, diversity-aware strategies to ensure reliable misinformation detection. DriftBench provides a foundation for advancing trustworthy verification methods in an increasingly complex and manipulated information landscape.

### v2

@@ Abstract
Introduction
Multimodal misinformation, often in the form of image-text combinations, poses escalating threats to public discourse, societal trust, and civic stability (bovet2019influence; murayama2021dataset; wang2024harmfully; deceptiondecoder). To counter this threat, Large Vision-Language Models (LVLMs) (DBLP:journals/corr/gpt-4o; DBLP:journals/corr/Gemini; llava; meta2024llama) has emerged as a dominant approach for multimodal misinformation detection (MMD), offering strong multimodal reasoning and retrieval-augmented verification capabilities (kangur2025multireflect; qi2024sniffer; braun2024defame; li2025imrrf; geng2024multimodal; wu2025beyond).
@@ Introduction
However, the rise of generative AI (GenAI) systems such as GPT-4o (DBLP:journals/corr/gpt-4o) is reshaping the information landscape. GenAI enables the creation of highly diverse and stylistically varied content at scale (kiskola2025generative; davenport2023generative; nishal2024envisioning; kieslich2024anticipating), fundamentally transforming how news is composed, visualized, and disseminated. We refer to this phenomenon as GenAI-driven News Diversity, which manifests in two key forms: (1) Controlled News Diversity, where users produce semantically consistent variants of the same news content through rewording, paraphrasing, or shifts in visual framing; and (2) Open-ended News Diversity, where GenAI generates entirely novel content via text-to-image synthesis or narrative reconstruction.
While this diversity enhances expressiveness and audience reach, it also introduces substantial challenges for LVLM-based MMD systems. Most existing models either assess image-text coherence using internal knowledge (wang2024mfc) or retrieve external evidence for fact verification (li2025cmie; qi2024sniffer). Both approaches become fragile under content variation, which gives rise to what we term multi-level drift. As illustrated in Figure 1, we identify two key drift phenomena: (1) Model-Level Misperception Drift: Variations in surface form, including lexical, syntactic, or visual aspects, may distort the model’s perception, even when semantics remain unchanged. This issue is exacerbated by LVLMs’ known hallucination tendencies (bai2024hallucination; liu2024hallucination). (2) Evidence-Level Drift: Content diversity compromises evidence retrieval by reducing the likelihood of finding exact matches, leading to loosely relevant or misleading external evidence (chung2023guard; xiao2021you; feng2020adversarial). Beyond natural drift, GenAI also enables adversarial manipulation through guided retrieval attacks, wherein malicious actors flood the web with strategically crafted synthetic content. Such contamination can bias retrieval pipelines, poisoning the fact-checking process and degrading LVLM performance (xu2024invisible; dai2023llms). Despite their growing risk, these vulnerabilities remain insufficiently studied.
@@
To this end, we introduce DriftBench, a large-scale benchmark consisting of 16,000 news instances spanning six diversified categories across both image and text modalities, with human validation to ensure semantic fidelity. Specifically, we source high-quality real and fake samples from NewsCLIPpings (newsclippings) and systematically diversify them using an automated GenAI-based pipeline grounded in our news diversity taxonomy. For real news, we apply controlled diversification to the image, text, or both while preserving factual consistency. For fake news, we incorporate both controlled and open-ended transformations, rewriting text to fabricate false narratives and generating misleading images aligned with the fabricated content. The resulting dataset includes eight structured categories, consisting of six diversified variants and two original classes, thereby enabling rigorous evaluation of LVLM robustness under GenAI-driven content variation.
With DriftBench, we design three corresponding evaluation tasks: (1) Performance under Diversity: Evaluate how SOTA LVLM-based detectors handle controlled and open-ended content variation, focusing on degradation due to model- and evidence-level drift. (2) Robustness to Malicious Evidence Contamination: Simulate guided retrieval attacks by injecting GenAI-generated misleading evidence into the verification pipeline. (3) Reasoning Behavior Analysis: Employ an LVLM-as-a-Judge framework (li2024judge-1; chen2024judge-2; pu2025judge-3) to assess models’ explanation quality and attribution patterns under diverse input conditions.
@@
NewsCLIPpings (newsclippings) OOC misinformation - - -
VERITE (papadopoulos2024verite) Real-world misinformation - - -
MMFakeBench (liu2024mmfakebench) Multi-source misinformation - - -
MFC-Bench (wang2024mfc) Multimodal fact-checking - - - -
MiRAGenNews (MiRAGeNews) AI-generated misinformation - - -
@@
LVLM-Based MMD Approaches
@@
LVLM-based multimodal fact-checking is the current mainstream method for multimodal misinformation detection. SNIFFER (qi2024sniffer) performs two-stage fine-tuning based on InstructBLIP (dai2023instructblip), aiming to detect misinformation from the perspective of entity matching and evidence verification. CMIE (li2025cmie) and E2LVLM (wu2025e2lvlm) further explore how to discover deeper relationships and optimize the selection and usage of external evidence. LRQ-FACT (lrq-fact) and LEMMA (xuan2024lemma) optimize retrieval strategies to acquire more targeted external evidence. These methods fully investigate the capabilities of LVLMs in retrieval, comprehension, and reasoning, and are optimized around the acquisition and utilization of high-quality evidence. However, most of them assume that the input claim is static and expressed in a single form, lacking systematic investigation and evaluation of such methods’ performance under news diversity and the resulting multi-level drift.
Multimodal Misinformation Benchmarks
@@
To evaluate multimodal misinformation detection (MMD), several benchmarks have been introduced. DGM4 (shao2023dgm4) targets fine-grained visual and textual manipulations. NewsCLIPpings (newsclippings) constructs out-of-context (OOC) cases by mismatching images and text with external evidence. VERITE (papadopoulos2024verite) collects real-world OOC misinformation from fact-checking platforms to analyze modality bias. MiRAGenNews (MiRAGeNews) employs GenAI techniques to synthesize misinformation image–text pairs, while MMFakeBench (liu2024mmfakebench) aggregates misinformation from both real and GenAI sources to assess robustness under mixed conditions. MFC-Bench (wang2024mfc) probes LVLMs’ internal fact-checking capabilities.
@@
In contrast, DriftBench is the first benchmark to systematically examine challenges introduced by GenAI-driven news diversity, with a focus on two overlooked forms of drift: model misperception drift and evidence drift. By generating semantically diverse real and fabricated variants through both controlled and open-ended transformations, it addresses a critical gap in existing benchmarks. A comparative overview is provided in Table 1.
DriftBench
@@
To systematically evaluate how GenAI-driven news diversity induces multi-level drift in LVLM-based multimodal misinformation detection, we introduce DriftBench, a large-scale benchmark of 16,000 news instances across eight categories. These categories comprise two base types (real and fake news with human-written text and authentic images) and six diversification strategies. Built through an automated GenAI-based synthesis pipeline applied to real-world news, DriftBench enables scalable, controlled, and semantically aligned diversification. Each instance contains an image–text pair—either original or diversified—together with externally retrieved web evidence. An overview of the dataset construction and evaluation setup is shown in Figure 2. Our code, data, and supplementary materials are publicly available [^1] https://github.com/fanxiao15/DriftBench.
@@
Taxonomy of GenAI-Driven News Diversity
@@
  • F_OI_OT: An out-of-context (OOC) misinformation instance (newsclippings; qi2024sniffer), created by mismatching the image and text from two different but individually trustworthy news reports—misleading when paired together.
@@
Dataset Creation
Based on the proposed taxonomy, we construct DriftBench, a large-scale benchmark comprising real and fake news instances systematically diversified through GenAI-driven content variation. We begin by sourcing 4,000 high-quality samples (2,000 real and 2,000 fake) from the NewsCLIPpings dataset (newsclippings). Real instances comprise semantically aligned image-caption pairs published by reputable news outlets. Fake instances are constructed via OOC manipulation, where the image and text are drawn from two different trustworthy reports. While individually accurate, their combination results in a misleading narrative with high surface plausibility.
@@
Text diversifications and fabrications are generated using GPT-4o (DBLP:journals/corr/gpt-4o). For text diversification (DT), we prompt the model to rephrase the original text while preserving its semantics. For text fabrication (FT), we prompt the model to generate a false narrative by altering the described event. Image diversifications (DI) are generated using FLUX.1 Redux [dev] (an open-source image variant generation model) (labs2025flux1kontextflowmatching), the model generates visually varied but semantically consistent depictions of the original content. Image fabrications (FI) are generated using FLUX.1 [dev] (an open-source text-to-image model) (flux2024), we condition the image generation on the fabricated narrative FT, yielding fully synthetic misinformation instances.
@@
For each image-text pair, we retrieve external evidence to support LVLM-based multimodal fact-checking. For image-based retrieval, we use a web crawler to identify webpages containing the image and extract their page titles as evidence. For text-based retrieval, we adopt the query generation strategy from LEMMA (xuan2024lemma), using the Serper API to retrieve relevant web documents.
@@
Data Quality Assessment
@@
We report two metrics to assess data quality: (1) semantic preservation accuracy, computed from aggregated human annotations, and (2) inter-annotator agreement using Fleiss’ $\kappa$ (fleiss1971measuring). Results show that DriftBench exhibits high semantic consistency and labeling reliability, with 98.6% accuracy and a Fleiss’ $\kappa$ of 0.872 for real-instance variants, and 96.6% accuracy with a Fleiss’ $\kappa$ of 0.741 for controlled fake variants.
Evaluation Tasks and Metrics
@@
Following established practices in LVLM-based MMD evaluation (li2025cmie; xuan2024lemma; qi2024sniffer), we report Accuracy and F1 as the primary metrics for detection performance.
Experiments
Experimental Setup
@@
Vanilla LVLMs. We assess three widely used models: (1) GPT-4o-mini (DBLP:journals/corr/gpt-4o): a proprietary large vision-language model. (2) Claude-3.7-Sonnet (claude3_7): a large reasoning model with multimodal capabilities. (3) Qwen-VL-Plus (bai2023qwen-vl): a strong open-source vision-language model. Prompting strategies and instruction templates used for Vanilla LVLMs are detailed in Appendix (Figure 11).
LVLM-Based MMD Methods. We also evaluate three state-of-the-art MMD approaches. (1) SNIFFER (qi2024sniffer), which fine-tunes LVLMs and retrieves external evidence directly through image-based querying. (2) CMIE (li2025cmie), which models deep semantic alignment between image and text while mitigating noise in retrieved evidence. and (3) LEMMA (xuan2024lemma), which employs a modular pipeline that first generates optimized queries and then retrieves high-quality evidence.
Performance under GenAI-Driven Diversity
@@
Robustness to Evidence Contamination
Contaminated evidence further degrades model performance and can steer predictions in misleading directions. As discussed in Section Evaluation Tasks and Metrics, malicious actors can manipulate retrieval pipelines by injecting fabricated or misleading content into the evidence set. Table 3 shows that introducing such adversarial evidence causes a clear performance drop across all evaluated methods. Detailed results in Appendix A.6 indicate that real instances are increasingly misclassified as fake when contaminated evidence is present. For fake instances, the misleading evidence often supports the false narrative, leading models to mistakenly classify them as real and thus producing the outcome intended by the adversary.
@@
Table 3: Comparison of LVLM-based multimodal misinformation detection under malicious evidence contamination, evaluated on the DI_OT setting with controlled news diversity.
@@
Model Reasoning Behavior Analysis
Figure 5: Reasoning explanations generated by LVLMs under controlled settings (R_OI_OT, R_DI_OT), open ended scenarios (F_FI_FT), and cases involving malicious evidence contamination (F_DI_OT).
@@
Figure 6: Error attribution showing the distribution of error types across different methods for real and fake instances.
@@
Conclusion
We present DriftBench, the first benchmark explicitly designed to evaluate the vulnerabilities of LVLM-based misinformation detection in the face of GenAI-driven news diversity. Our diversified benchmark demonstrates that both controlled and open-ended content variation, along with adversarial evidence contamination, can substantially weaken the performance of current LVLM-based MMD systems. Beyond accuracy degradation, our analysis reveals distinct reasoning behaviors and error patterns among different model classes, reflecting varied reliance on visual-textual coherence and external evidence. These findings emphasize the need for robust, diversity-aware strategies to ensure reliable misinformation detection. DriftBench provides a foundation for advancing trustworthy verification methods in an increasingly complex and manipulated information landscape.
Acknowledgment
This work is supported by the Yunnan Province expert workstations (Grant No. 202305AF150078), National Natural Science Foundation of China (Grant Nos. 62162067, 62562061, 62502422 and 62462067), Yunnan Fundamental Research Project (Grant Nos. 202401AT070474, 202501AU070059), Yunnan Province Special Project (Grant No.202403AP140021), Yunnan Provincial Department of Education Science Research Project (Grant Nos. 2025J0006, 2024J0010 and 2025J0007), Scientific Research and Innovation Project of Postgraduate Students in the Academic Degree of Yunnan University (KC-4248590) and China Scholarship Council (CSC) program. This research is also supported by the Ministry of Education, Singapore, under its MOE AcRF TIER 3 Grant (MOE-MOET32022-0001).